OpenStudy (anonymous):

Solve: √ 9x + 10 = x

OpenStudy (saifoo.khan):

\[\sqrt{9x+10} = x\]?

OpenStudy (anonymous):

lol same question as him. but if its sqrt(9x)+10=x then x=25 but if it is how saifoo.khan wrote it than it is x=10